OpenStudy (anonymous):

The number of quadratic equations having real roots and which do not change by squaring their roots? (a)4 (b)3 (c)2 (d)1

OpenStudy (anonymous):

D

OpenStudy (anonymous):

can u plz explain how ???

OpenStudy (anonymous):

1^0.5=1^2=1